OpenStudy (anonymous):

Use the Subtraction Property of Equality to complete the following statement: if 10x+6=21 then ___=15 a)10x – 6 b)10x c)10x + 6 d)1.5

OpenStudy (blurbendy):

What happens when you subtract 6 from both sides? What does the equation become?

OpenStudy (blurbendy):

I'm referring to the first equation, 10x + 6 = 21

OpenStudy (anonymous):

you get d

OpenStudy (anonymous):

mind helping me on another on??

OpenStudy (blurbendy):

10x + 6 = 21 -6 -6 10x = 15 I'm not sure if the question is asking you to solve for x though

OpenStudy (anonymous):

Which property justifies the following statement? If x=4 and y=3x-11 then y=3x4-11 A)substitution property of equality. b)subtraction property of equality c)division property of equality D)addition property of equality i thought it was a but im not sure

OpenStudy (anonymous):

and it was i got it right ! thanks !

OpenStudy (blurbendy):

Okay, cool

OpenStudy (blurbendy):

then y = 3(4) - 11 ? Is that what the last part says?

OpenStudy (anonymous):

yupp

OpenStudy (blurbendy):

Okay, since we knew x = 4, all we did was "substitute" that in for x in the other equation.

OpenStudy (anonymous):

ok so i was right it was a ! thanks !

OpenStudy (blurbendy):

yw!